# Unlock the Power of Algorithm Features: A Generalization Analysis for Algorithm Selection

### v1

@@ 3 Generalization Error of the Adaptive Feature-based Model
Follow the same definition in Theorem 2: (1) Let $L_{r}$ and $\gamma_{r}$ be the Lipschitz constants of the model MLPr and its loss function. $c=c_{0}(1+\eta)+\sqrt{\frac{1}{2}(1+\eta)\ln\frac{1}{\delta}}$. $Error_{\mathcal{T}}(\text{MLP}_{r})$ and $Error_{\mathcal{S}}(\text{MLP}_{r})$ denote the test error and empirical error of Model MLPr, then with probability of at least $1-\delta$,
$Error_{\mathcal{T}}(\text{MLP}_{r})\leq Error_{\mathcal{S}}(\text{MLP}_{r})+$ (24)
$\mathcal{O}\left(\frac{\gamma_{r}L_{r}(1+\eta)}{\eta|\mathcal{S}_{\mathcal{P}}|}\right)+\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{P}}|}}\right)$
(2) Let $L_{m}$ and $\gamma_{m}$ be the Lipschitz constants of the model MLPm and its loss function. $c=c_{0}(1+\eta)+\sqrt{\frac{1}{2}(1+\eta)\ln\frac{1}{\delta}}$. $Error_{\mathcal{T}}(\text{MLP}_{m})$ and $Error_{\mathcal{S}}(\text{MLP}_{m})$ denote the test error and empirical error of Model MLPm, then with probability of at least $1-\delta$,
$Error_{\mathcal{T}}(\text{MLP}_{m})\leq Error_{\mathcal{S}}(\text{MLP}_{m})+$ (25)
$\mathcal{O}\left(\frac{\gamma_{m}L_{m}|\mathcal{S}_{\mathcal{A}}|(1+\eta)}{\eta|\mathcal{S}_{\mathcal{P}}|}\right)+\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{P}}|}}\right)$
@@
The proof can be followed by Theorem 2, where the proof process of model MLPm further involves the application of Lemma 1 in [38], to solve the multi-class scenario. ∎
Herein, we focus on analyzing the relationship between the upper bounds of generalization error for different models and the size of the training instances. According to Corollary 2, we find that, in the case of a finite training instance size, the generalization error upper bound of MLPr is represented by a slack term of $\mathcal{O}\left(\frac{\gamma_{r}L_{r}(1+\eta)}{\eta|\mathcal{S}_{\mathcal{P}}|}\right)$ in relation to the training set. This indicates a weaker generalization error capability compared to Modela, which exhibits a slack term of $\mathcal{O}\left(\frac{\gamma L\|W^{(0)}\|_{2}(1+\eta)}{\eta|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}\right)$. Moreover, the generalization error of MLPm is even proportional to the number of candidate algorithms, following the slack term $\mathcal{O}\left(\frac{\gamma_{m}L_{m}|\mathcal{S}_{\mathcal{A}}|(1+\eta)}{\eta|\mathcal{S}_{\mathcal{P}}|}\right)$. These results underscore the advantages of utilizing adaptive algorithm features, particularly when there is a relatively large number of candidate algorithms. The incorporation of adaptive algorithmic features is beneficial for enhancing the model’s generalization performance. In the case of abundant training samples, adaptive algorithm features demonstrate the same advantages in generalization performance. The overall asymptotic rate of MLPr, MLPm, and Modela follow the slack term $\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{P}}|}}\right)$, $\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{P}}|}}\right)$, and $\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}}\right)$, respectively.
@@ 4 Generalization Error of the Predefined Feature-based Model
where $\mathcal{F}$ is a class of real-valued network, $\mathcal{N}_{i}^{j}$ denotes the sub-network from $i$-th layer to $j$-th layer in model $\mathbf{f}\left(x_{i}\right)$, $W^{(i)}$ denote the parameters of the $i$-th layer and $\phi^{(i)}(*)$ denote the $1$-Lipschitz and positive-homogeneous activation function. To analyze the upper bound of inductive Rademacher complexity $\mathfrak{R}_{|\mathcal{S}|}(\mathcal{F})$, we first provide two auxiliary lemmas in [39] and [40]:
@@
Lemma 3 will be used in inequality scaling in the proof of Theorem 3, and Lemma 4 will be used to bound the expectation term. Hence, we first introduce a parameter $\lambda$ and the calculation $\exp$, to construct the form in Eq. (28), as shown in the proof of Theorem 3 as follows.
@@
According to Jensen’s inequality for concave functions and Cauchy-Schwarz inequality, ^$\hat{\mathfrak{R}}_{|\mathcal{S}|}\left(\mathcal{F}\right)$ satisfies:
@@
According to Theorem 4, we can observe that the inductive generalization error is related to three factors: the training scale $|\mathcal{S}|$, the model parameter-related factor $\Gamma_{\mathbf{f}}$ and loss function parameter $\gamma$, and the value of training data $\Gamma_{\mathcal{S}}$. By Eq. (39), in most cases where $\Gamma_{\mathbf{f}},\Gamma_{\mathcal{S}}>1$ and $\sqrt{|\mathcal{S}|}\gg(\log 2)l$, Theorem 4 can be simplified to the conclusion in Corollary 3, which states that the rate of change is governed by $\mathcal{O}\left(\frac{\Gamma_{\mathbf{f}}\Gamma_{\mathcal{S}}^{\frac{3}{4}}}{|\mathcal{S}_{\mathcal{A}}|^{\frac{1}{4}}|\mathcal{S}_{\mathcal{P}}|^{\frac{1}{4}}\gamma}\right)$. In rare cases where $\Gamma_{\mathbf{f}},\Gamma_{\mathcal{S}}<1$ or $\sqrt{|\mathcal{S}|}\approx(\log 2)l$, the rate of change in generalization error becomes $\mathcal{O}\left(\frac{\sqrt{(\log 2)l\Gamma_{\mathbf{f}}\Gamma_{\mathcal{S}}}}{\sqrt{|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}\gamma}\right)+\mathcal{O}\left(\frac{c_{1}}{\sqrt{|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}}\right)$. This means that in scenarios with fewer training samples and a relatively simple model, the improvement of model generalization by training samples is more significant.
@@
Different from the transductive settings, the generalization error in Theorem 4 is not affected by the size of the test set, which is a characteristic brought by inductive learning. Furthermore, in inductive manner, the increased number of candidate algorithms also has a positive impact on reducing generalization error, which is similar to the transductive generalization performance. According to the multiplication invariance of the $L_{2}$-norm in Eq. (12), both Theorem 2 and Theorem 4 are affected by the product of the norms of each layer’s parameter matrix, while the generalization performance of intuitive learning is influenced by the Frobenius norm, and the generalization performance of transductive learning is influenced by the $L_{2}$-norm. The most significant difference between inductive and transductive settings, is the impact of training instances scale, at a rate of $\frac{1}{|\mathcal{S}_{\mathcal{P}}|^{\frac{1}{4}}|\mathcal{S}_{\mathcal{A}}|^{\frac{1}{4}}}$ in inductive settings.
In conclusion, within the inductive setting, we observe a more pronounced difference on relationship between generalization error and training instances scale compared to transductive settings, at a rate of $\frac{1}{|\mathcal{S}_{\mathcal{P}}|^{\frac{1}{4}}|\mathcal{S}_{\mathcal{A}}|^{\frac{1}{4}}}$. However, due to their distinct generalization implications, these two boundaries should not be directly comparable. The superior asymptotic rate in transductive learning arises from the assumption that the test and training sets originate from the same distribution. We will illustrate in a corollary of Theorem 4 how distribution shift impacts generalization error in inductive learning.
@@
Since the generalization error in the inductive setting pertains specifically to the utilization of predefined features, the subsequent analysis primarily provides some practical advice for incorporating predefined features. Firstly, in terms of the impact of instance value on the generalization performance, adaptive features are superior to predefined features because adaptive features are learned by the model and their values are more stable. Therefore, the influence of $\|W^{(0)}\|_{2}$ on the generalization error in Theorem 2 is limited, while the value of $\Gamma_{\mathcal{S}}$ in Theorem 4 is unpredictable. This indicates that the generalization performance of the model under predefined features should consider the stability and value range of the predefined features. In practical applications, these factors should be considered in the selection of predefined models and the preprocessing of features. Compared to using adaptive features, another advantage of predefined features is their generalization on new algorithms. Below, we simultaneously consider the distribution shift between training and test sets, as well as the impact of negative sampling, to obtain a corollary about generalization from Theorem 4.
@@
Follow the same definition in Theorem 4, and $\mathcal{T}_{\mathcal{A}}-\mathcal{S}_{\mathcal{A}}\neq\varnothing$. Let $\chi^{2}(P_{\mathcal{T}}\|P_{\mathcal{S}})=\int\frac{P^{2}_{\mathcal{T}}(x_{j})}{P_{\mathcal{S}}(x_{j})}-1$ denote the chi-square divergence. If $\prod_{i=1}^{l}R_{i}>1$ and $\sqrt{|\mathcal{S}_{\mathcal{P}}|\cdot|\mathcal{S}_{\mathcal{A}}|}\gg(\log 2)l$, then with probability at least $1-\delta$ over the sample, we have,
@@
Compared to adaptive features, predefined features demonstrate generalization on new algorithms that are not present in the training data. Corollary 4 explains the impact of distribution shift between training and test data on generalization error. It can be observed that as the chi-square divergence between the distributions increases, the generalization error gradually increases at a rate of $[\chi^{2}(P_{\mathcal{T}}\|P_{\mathcal{S}})+1]^{\frac{3}{4}}$. Similarly, the size of the problems and candidate algorithms in the training data still affect the generalization error at a rate of $\frac{1}{|\mathcal{S}_{\mathcal{P}}|^{\frac{1}{4}}|\mathcal{S}_{\mathcal{A}}|^{\frac{1}{4}}}$. In addition to the impact of sample size, it has been noted that in scenarios involving distribution shift, the complexity of a model exacerbates generalization error. Thus, when handling distribution shift in practical applications, especially when aiming for generalization capacity on candidate algorithms, it is advisable to refrain from employing excessively complex models.
@@ 5 Experiment
The mathematical bounds, though often too loose for practical applications, can provide a rough estimate of the sample size required for a certain level of performance. This section presents experimental results aimed at validating some of the theoretical results proposed in this paper. In order to examine the influence of changes in distribution and quantity on model performance, it is necessary to adjust the distribution of training and testing data during the theoretical validation process. Hence, the experiments in this study are conducted on simulated data, which provided a controlled setting where the exact underlying distributions of problems and algorithms could be known. In the subsequent subsection, we introduce the methodology employed for data simulation.
@@ 6 Conclusion
Algorithm selection is a crucial step in automated machine learning. Currently, most mainstream techniques rely on problem features for modeling, while literature discussing algorithm features is relatively scarce. Despite a few empirical studies that have provided evidence of the effectiveness of algorithm features, the potential benefits of incorporating algorithm features into algorithm selection models and their suitability for different scenarios remain unclear. This ambiguity arises from the fact that while algorithm features can provide additional information to enhance the model’s generalization under distribution shift, especially when faced with new candidate algorithms, the complexity introduced by handling algorithm features often requires a larger training dataset. Consequently, algorithm feature-based models are influenced by various factors, and empirical studies alone cannot fully elucidate the underlying mechanisms behind performance variations.
In this study, we addressed the existing gap in algorithm selection research by proposing a provable guarantee for algorithm feature-based algorithm selection model and taking a generalization perspective. We provided a solid foundation for understanding the mechanisms behind algorithm selection models. Our analysis encompassed adaptive algorithm features and predefined algorithm features under different learning settings, and we derived upper bounds for the generalization error of models based on each type of feature using the Rademacher complexity. Our theoretical investigations provide valuable insights into the practical application of algorithm features in real-world scenarios. Furthermore, our experimental results demonstrated the superiority of algorithm feature-based models, particularly in complex multi-algorithm problems, showcasing superior generalization capabilities. These algorithm feature-based models also proved to be well-suited for deployment in scenarios characterized by distribution shifts, where the generalization error exhibited a proportional relationship with the chi-square distance between the training and test sets. Overall, our study highlights the importance of considering algorithm features in algorithm selection and provides a solid foundation for future research in this area. By incorporating algorithm features, researchers and practitioners can harness the benefits of enhanced generalization and make informed decisions when selecting algorithms for various scenarios.

### v2

@@ 3 Generalization Error of the Adaptive Feature-based Model
Follow the same definition in Theorem 2: (1) Let $L_{r}$ and $\gamma_{r}$ be the Lipschitz constants of the Modelreg and its loss function. $c=c_{0}(1+\eta)+\sqrt{\frac{1}{2}(1+\eta)\ln\frac{1}{\delta}}$. $Error_{\mathcal{T}}(\text{Model}_{reg})$ and $Error_{\mathcal{S}}(\text{Model}_{reg})$ denote the test error and empirical error of Modelreg, then with probability of at least $1-\delta$,
$Error_{\mathcal{T}}(\text{Model}_{reg})\leq Error_{\mathcal{S}}(\text{Model}_{reg})+$ (24)
$\mathcal{O}\left(\frac{\gamma_{r}L_{r}\|W^{(0)}\|_{2}(1+\eta)}{\eta|\mathcal{S}_{\mathcal{P}}|}\right)+\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{P}}|}}\right)$
(2) Let $L_{c}$ and $\gamma_{c}$ be the Lipschitz constants of the Modelcla and its loss function. $c=c_{0}(1+\eta)+\sqrt{\frac{1}{2}(1+\eta)\ln\frac{1}{\delta}}$. $Error_{\mathcal{T}}(\text{Model}_{cla})$ and $Error_{\mathcal{S}}(\text{Model}_{cla})$ denote the test error and empirical error of Modelcla, then with probability of at least $1-\delta$,
$Error_{\mathcal{T}}(\text{Model}_{cla})\leq Error_{\mathcal{S}}(\text{Model}_{cla})+$ (25)
$\mathcal{O}\left(\frac{\gamma_{m}L_{m}|\mathcal{S}_{\mathcal{A}}|\|W^{(0)}\|_{2}(1+\eta)}{\eta|\mathcal{S}_{\mathcal{P}}|}\right)+\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{P}}|}}\right)$
@@
The proof can be followed by Theorem 2, where the proof process of the Modelcla further involves the application of Lemma 1 in [38], to solve the multi-class scenario. ∎
Herein, we focus on analyzing the relationship between the upper bounds of generalization error for different models and the size of the training instances. According to Corollary 2, we find that, in the case of a finite training instance size, the generalization error upper bound of Modelreg is represented by a slack term of $\mathcal{O}\left(\frac{\gamma_{r}L_{r}\|W^{(0)}\|_{2}(1+\eta)}{\eta|\mathcal{S}_{\mathcal{P}}|}\right)$ in relation to the training set. This indicates a weaker generalization capability in terms of the growing number of algorithms, compared to Modela which exhibits a slack term of $\mathcal{O}\left(\frac{\gamma L\|W^{(0)}\|_{2}(1+\eta)}{\eta|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}\right)$. Moreover, the upper bound of Modelcla’s generalization error is even proportional to the number of candidate algorithms, following the slack term $\mathcal{O}\left(\frac{\gamma_{c}L_{c}\|W^{(0)}\|_{2}|\mathcal{S}_{\mathcal{A}}|(1+\eta)}{\eta|\mathcal{S}_{\mathcal{P}}|}\right)$. These results underscore the advantages of utilizing adaptive algorithm features, particularly when there is a relatively large number of candidate algorithms. The incorporation of adaptive algorithmic features is beneficial for enhancing the model’s generalization performance. In the case of abundant training samples, adaptive algorithm features demonstrate the same advantages in generalization performance. The overall asymptotic rate of Modelreg, Modelcla, and Modela follow the slack term $\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{P}}|}}\right)$, $\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{P}}|}}\right)$, and $\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}}\right)$, respectively.
@@ 4 Generalization Error of the Predefined Feature-based Model
where $\mathcal{F}$ is a class of real-valued network, $\mathcal{N}_{i}^{j}$ denotes the sub-network from $i$-th layer to $j$-th layer in model $\mathbf{f}\left(x_{i}\right)$, $W^{(i)}$ denotes the parameters of the $i$-th layer and $\phi^{(i)}(*)$ denotes the $1$-Lipschitz and positive-homogeneous activation function. To analyze the upper bound of inductive Rademacher complexity $\mathfrak{R}_{|\mathcal{S}|}(\mathcal{F})$, we first provide two auxiliary lemmas in [39] and [40]:
@@
Lemma 3 will be used in inequality scaling in the proof of Theorem 3, and Lemma 4 will be used to bound the expectation term. Hence, we first introduce a parameter $\lambda$ and the calculation ‘$\exp$’, to construct the form in Eq. (28), as shown in the proof of Theorem 3 as follows.
@@
According to Jensen’s inequality for concave functions and Cauchy-Schwarz inequality, ^$\hat{\mathfrak{R}}_{|\mathcal{S}|}\left(\mathcal{F}\right)$ satisfies the following inequality:
@@
According to Theorem 4, we can observe that the inductive generalization error is related to three factors: the training scale $|\mathcal{S}|$, the model parameter-related factor $\Gamma_{\mathbf{f}}$ and loss function parameter $\gamma$, and the value of training data $\Gamma_{\mathcal{S}}$. By Eq. (39), in most cases where $\Gamma_{\mathbf{f}},\Gamma_{\mathcal{S}}>1$ and $\sqrt{|\mathcal{S}|}\gg(\log 2)l$, Theorem 4 can be simplified to the conclusion in Corollary 3, which states that the rate of change is governed by $\mathcal{O}\left(\frac{\Gamma_{\mathbf{f}}\Gamma_{\mathcal{S}}^{\frac{3}{4}}}{|\mathcal{S}_{\mathcal{A}}|^{\frac{1}{4}}|\mathcal{S}_{\mathcal{P}}|^{\frac{1}{4}}\gamma}\right)$. In rare cases where $\Gamma_{\mathbf{f}},\Gamma_{\mathcal{S}}<1$ or $\sqrt{|\mathcal{S}|}\approx(\log 2)l$, the rate of change in generalization error becomes $\mathcal{O}\left(\frac{\sqrt{(\log 2)l\Gamma_{\mathbf{f}}\Gamma_{\mathcal{S}}}}{\sqrt{|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}\gamma}\right)+\mathcal{O}\left(\frac{c_{1}}{\sqrt{|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}}\right)$. This means that in scenarios with fewer training instances and a relatively simple model, the improvement of model generalization by training samples is more significant.
@@
Different from the transductive settings, the generalization error in Theorem 4 is not affected by the size of the test set, which is a characteristic brought by inductive learning paradigm. Furthermore, in inductive manner, the increased number of candidate algorithms also has a positive impact on reducing generalization error, which is similar to the transductive generalization performance. According to the multiplication invariance of the $L_{2}$-norm in Eq. (12), both Theorem 2 and Theorem 4 are affected by the product of the norms of each layer’s parameter matrix, while the generalization performance of intuitive learning is influenced by the Frobenius norm, and the generalization performance of transductive learning is influenced by the $L_{2}$-norm. Additionally, within the inductive setting, we observe a more pronounced difference on relationship between generalization ability and training instances scale compared to transductive settings, at a rate of $\frac{1}{|\mathcal{S}_{\mathcal{P}}|^{\frac{1}{4}}|\mathcal{S}_{\mathcal{A}}|^{\frac{1}{4}}}$. However, due to their distinct generalization implications, these two boundaries should not be directly comparable. The superior asymptotic rate in transductive learning arises from the assumption that the test and training sets originate from the same distribution. We will illustrate in a corollary of Theorem 4 how distribution shift impacts generalization error in inductive learning.
@@
Since the generalization error in the inductive setting pertains specifically to the utilization of predefined features, the subsequent analysis primarily provides some practical advice for incorporating predefined features. Firstly, in terms of the impact of instance value on the generalization performance, adaptive features are superior to predefined features because adaptive features are learned by the model and their values are more stable. Therefore, the influence of $\|W^{(0)}\|_{2}$ on the generalization error in Theorem 2 is limited, while the value of $\Gamma_{\mathcal{S}}$ in Theorem 4 is unpredictable. This indicates that the generalization performance of the model under predefined features should consider the stability and value range of the predefined features. Take large language model-extracted features as an example. In practical applications, these factors should be considered in the selection of pretrained models and the preprocessing of features. Compared to using adaptive features, another advantage of predefined features is their generalization on new algorithms. Below, we simultaneously consider the distribution shift between training and test sets, as well as the impact of negative sampling, to obtain a corollary about generalization from Theorem 4.
@@
Follow the same definition in Theorem 4, and $\mathcal{T}_{\mathcal{A}}-\mathcal{S}_{\mathcal{A}}\neq\varnothing$. Let $\chi^{2}(P_{\mathcal{T}}\|P_{\mathcal{S}})=\int\frac{P^{2}_{\mathcal{T}}(x_{j})}{P_{\mathcal{S}}(x_{j})}-1$ denote the chi-square divergence between training and test distributions. If $\prod_{i=1}^{l}R_{i}>1$ and $\sqrt{|\mathcal{S}_{\mathcal{P}}|\cdot|\mathcal{S}_{\mathcal{A}}|}\gg(\log 2)l$, then with probability at least $1-\delta$ over the sample, we have,
@@
Compared to adaptive features, predefined features demonstrate generalization on new algorithms that are not present in the training data. Corollary 4 explains the impact of distribution shift between training and test data on upper bound of the generalization error. It can be observed that as the chi-square divergence between the distributions increases, the upper bound of the generalization error gradually increases at a rate of $[\chi^{2}(P_{\mathcal{T}}\|P_{\mathcal{S}})+1]^{\frac{3}{4}}$. Similarly, the size of the problems and candidate algorithms in the training data still affect the generalization error at a rate of $\frac{1}{|\mathcal{S}_{\mathcal{P}}|^{\frac{1}{4}}|\mathcal{S}_{\mathcal{A}}|^{\frac{1}{4}}}$. In addition to the impact of sample size, it has been noted that in scenarios involving distribution shift, the complexity of a model exacerbates generalization error. Thus, when handling distribution shift in practical applications, especially when aiming for generalization capacity on candidate algorithms, it is advisable to refrain from employing excessively complex models.
@@ 5 Experiment
The mathematical bounds, though often too loose for practical applications, can provide a rough estimate of the sample size required for a certain level of performance. This section presents experimental results aimed at validating some of the theoretical results proposed in this paper. In order to examine the influence of changes in distribution and quantity on model performance, it is necessary to adjust the distribution of training and testing data during the theoretical validation process. Hence, the experiments in this study are conducted on simulated data, which provided a controlled setting where the exact underlying distributions of problems and algorithms could be known. In Section 5.1, we describe the process of generating simulated data for continuous optimization problems. Then, Sections 5.2 and 5.3 investigate how varying the number of training problem instances and candidate algorithms affects model performance, respectively. In Section 5.4, we explore model performance under distribution shifts in training and testing data, and in Section 5.5, we analyze how increasing training sample size influences model performance under different distribution shifts. Finally, Section 5.6 looks at the relationship between model complexity and performance, especially under distribution shifts, to determine the optimal model size.
@@ 5.4 Impact of the Distribution Shift
Figure 4: The impact of the distribution shift on model performance.
@@ 5.6 Impact of the Model Complexity
Figure 6: The impact of the model complexity on model performance.
@@ 6 Conclusion
This paper delves into the complex interplay between algorithm features and problem features in the realm of algorithm selection, offering both theoretical and empirical insights. Our comprehensive study reveals several key findings and provides actionable guidelines for the practical application of algorithm features in real-world scenarios.
Our theoretical analysis highlights the distinct impacts of training set size on the Rademacher complexity under both transductive and inductive learning paradigms. We found that transductive learning models benefit more from larger training sets in terms of reducing Rademacher complexity compared to inductive learning models. This insight underscores the importance of tailoring training strategies to the learning paradigm employed. In the context of transductive learning, our results indicate that algorithm features improve model generalization, especially when dealing with a large number of candidate algorithms. The generalization error bounds for models leveraging algorithm features decrease more rapidly with increasing training data size, compared to models based solely on problem features. This advantage becomes more pronounced as the amount of training data grows, suggesting that algorithm feature-based models are particularly suited for scenarios with extensive training problems and a large number of candidate algorithms. When considering distribution shifts between the training and test sets, models based on algorithm features generally exhibit better generalization than those relying only on problem features. However, predefined algorithm features demonstrate superior generalization performance compared to adaptive features. While adaptive features adapt to the training data, they lack the robustness to handle new candidate algorithms in different distributions effectively. On the other hand, predefined features, which capture the intrinsic properties of algorithms, provide a more consistent generalization capability even under significant distribution shifts. Nonetheless, even predefined features suffer an increase in generalization error proportional to the chi-square distance between the training and test distributions. This finding emphasizes the need for careful consideration of distribution shifts in practical applications and suggests a cautious approach to using large-scale models in such scenarios.
In summary, our research provides a nuanced understanding of when and how to utilize algorithm features effectively. We recommend leveraging algorithm features in settings with a large number of candidate algorithms and considering the use of predefined features to handle potential distribution shifts. For situations involving significant distribution shifts, predefined features offer a more robust option, though the model complexity should be managed to mitigate potential increases in generalization error. By following these guidelines, practitioners can better harness the power of algorithm features to improve the accuracy and reliability of algorithm selection processes.